OpenStudy (anonymous):

(x-5)^2=81 my options are -5+-squareroot(81), 5+-squareroot(81), -25+-squareroot(81)

OpenStudy (anonymous):

\[-5\pm \sqrt{(81)}, 5\pm \sqrt{(81)}, -25\pm \sqrt{(81)}\]

OpenStudy (anonymous):